OpenStudy (anonymous):

HELP HELP HELP Find the sum of the three expressions and choose the correct answer. -8a^3b + 3a^2b^2 - 5 2 + 5a^2b^2 3a^3b + 7 - 9a^2b^2

OpenStudy (anonymous):

@dragonknight326 @hartnn @natedog12 @jacob1233

OpenStudy (anonymous):

just to let u know i am in 6th grade and i am 12 years old

OpenStudy (anonymous):

nevermind. sorry. highschool algebra here. you are going to enter living hell when you get to this.

OpenStudy (anonymous):

i know noooooooooooooooooooooooooooooooooo

OpenStudy (anonymous):

im currently taking my two final exams online and its gonna make or break my future. also I'm taking geometry next. then algebra 2, then in 12th grade i dont have to take a math subject so im just waiting for that day

OpenStudy (anonymous):

and i suck at everything math. once you add letters in there im as lost as a puppy

OpenStudy (anonymous):

my sister is a softmore

OpenStudy (anonymous):

sophmore* (grammar nazi sometimes)

OpenStudy (anonymous):

sophomore*

OpenStudy (jadeishere):

so it's based off combining like terms >.< -8a^3b + 3a^2b^2 - 5 2 + 5a^2b^2 3a^3b + 7 - 9a^2b^2 3a^2b^2, -9a^2b^2 and 5a^2b^2 are like terms so combine those and you get; -a^2b^2 then combine -8a^3b and 3a^3b so far -a^2b^2 - 5a^3b then there's 2 and 7 = 9 -a^2b^2 - 5a^3b + 9

OpenStudy (anonymous):

so the answer is...???

OpenStudy (anonymous):

these are the options. -5a^3b - a^2b^2 + 2 -5a^3b + 8a^2b^2 + 4 -5a^3b - a^2b^2 + 4

OpenStudy (jadeishere):

lol there's no answer to the expression. The simplified version is what i put above, just rearrange the terms >.<

OpenStudy (jadeishere):

Oh there's a -5, so -a^2b^2 - 5a^3b + 4

OpenStudy (anonymous):

LIFESAVER @Jadeishere thankyouthankyouthankyou

OpenStudy (anonymous):

can you help me some more? life on the line right now

OpenStudy (jadeishere):

Yeah :)

OpenStudy (anonymous):

Simplify: a - {4b - [a - (5b - c) + 3c - (a - b - 9c)]}

OpenStudy (jadeishere):

a - (4b - a - 5b + c + 3c - a + b + 9c) (PEMDAS) combine like terms within the parentheses 4b - 5b + b = -3b -a - a = -2a c + 3c + 9c = 13c a - (-3b - 2a + 13c) a + 3b + 2a - 13c 3b + 3a - 13c

OpenStudy (anonymous):

choices: -a - 8b + 13c a - 8b + 13c a - 2b + 13c -a - 2b + 13c

OpenStudy (jadeishere):

Let me look again..

OpenStudy (jadeishere):

a - 8b + 13c. This time i;m sure

OpenStudy (anonymous):

thanks. I have a bunch of problems, and right now math is the biggest one. need some serious help right now.

OpenStudy (jadeishere):

Kay >.<

OpenStudy (anonymous):

Choose the product. -6p^3(3p^2 + 5p - 1) -18p^5 - 30p^4 + 6p^3 -18p^6 - 30p^3 - 6p 18p^3 + 6p^2 - 30p^4 -18p^6 - 24p^3

OpenStudy (jadeishere):

first one

OpenStudy (anonymous):

Select the correct product. (x^2 + 6x + 9)(3x - 1) 3x^3 + 17x^2 + 21x - 9 3x^3 - 17x^2 - 21x - 9 3x^3+ 19x^2 + 27x + 9 3x^3+ 19x^2 + 9x - 9

OpenStudy (anonymous):

im just gonna say ultimate thank you's

OpenStudy (jadeishere):

3x^3 - x^2 + 18x^2 - 6x + 27x - 9 3x^3 + 17x^2 + 21x- 9

OpenStudy (anonymous):

Perform the indicated operation. (36m^2 - 60m + 25) ÷ (6m - 5) = _____ 6m + 5 6m - 5 R. 50 6m + 15 R. 100 6m - 5

OpenStudy (jadeishere):

6m - 5 >.<

OpenStudy (anonymous):

-x^2y^2 + x^4 + 4y^2 - 4x^2 (x - 2)(x - 2)(x + y)(x - y) (x + 2)(x + 2)(x + y)(x - y) (x + 2)(x - 2)(x + y)(x - y)

OpenStudy (anonymous):

^^^Factor the following polynomial completely.^^^